A sample of hard water contains calcium bicarbonate. 0.222g of Ca(OH)2 was used to soft and 50 ML of this sample.  Find out the amount of calcium bicarbonate per litre of the sample.
A) 9.72 gram/litre 
B) 5.34 gram/litre
C) 2.46 gram/litre
D) 4.35 gram/litre

Grade:11

1 Answers

Arun
3 years ago
Volume of the water = 50,000 liters.Now,10 liter of the water contains 1.62 grams of the Cacium bicarbonate.∴ 1 liter of the water contains 1.62/10 grams of the Calcium bicarbonate.∴ 50,000 litres of the water contains 0.162 × 50,000 = 8100 grams of the Calcium bicarbonate.When the lime water (Calcium Hydroxide) is added to the Calcium bicarbonate, it forms the Calcium carbonate and water, thus hardness of the water is removed. Reaction of its is ⇒ Ca(HCO₃)₂ + Ca(OH)₂ ------→ 2CaCO₃ + 2H₂OIn this Reaction we can see that the 1 mole of Lime water reacts with the one mole of Calcium Hydrogen Carbonate to remove the hardness of the water.∴ No. of moles of Calcium Hydrogen Bicarbonate = Mass/Molar Mass.   = 8100/162       = 50 molesAlso, No. of moles of Lime water = Mass/Molar mass  = Mass/74  Since no. of moles are equal, ∴ Mass/74 = 50⇒ Mass = 74 × 50∴ Mass =  3700 grams.Hence, the mass of the lime water required to remove the hardness of the water is 3700 grams.